# VisionLLaMA: A Unified LLaMA Interface for Vision Tasks

### v1

@@ 3 Augment (ours)
Table 14: Training procedures with ImageNet-1K.
B.2 Supervised Training of Pyramid VisionLLaMA
@@
B.3 Mask Image Modeling on ImageNet
@@
B.4 Linear Probing on ImageNet
@@
Table 15: Linear probing setting.
B.5 SFT on ImageNet for SSL-pre-trained Models
@@
Table 16: End-to-end fine-tuning setting for SSL.
B.6 Segmentation for SSL-pre-trained Models
@@
B.7 Segmentation for Pyramid Transformers
We follow the setting of [12], which is almost the same as B.6. We use a drop-path rate of 0.2 for the pyramid VisionLLaMA-B model.
B.8 Object Detection of Pyramid Transformers
[…]
B.9 Object Detection of Plain Transformers
[…]
B.10 Image Generation of DiT-LLaMA
We use the same VAE as [53]. We make use of the AdamW optimizer with momentum $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We use a global batch size of 256 across all models. The learning rate is fixed as 1 $\times$ $10^{-4}$. The training resolution is 256 $\times$ 256. As for inference, we use 250 steps DDPM. We keep the default setting of ADM [21] without tuning. Specifically, we use $t_{max}=1000$ linear schedule with $\beta$ from 0.0001 to 0.02 and learnable variance $\sigma_{\theta}$.
B.11 Image Generation of SiT-LLaMA
We use the same VAE as SD [53]. As for the ODE sampler, we utilize dopri5 and set atol and rtol to 1e-6 and 1e-3 respectively.
Appendix C Architecture Setting
C.1 Pyramid VisionLLaMA
@@
Table 17: Configuration details of Pyramid VisionLLaMA.
C.2 Plain Transformer for Vision Understanding
@@
C.3 Plain Transformer for Generation
@@
Table 19: Architecture settings for VisionLLaMA on image generation.

### v2

@@ 3 Augment
0.D.2 Supervised Training of Pyramid VisionLLaMA
@@
0.D.3 Mask Image Modeling on ImageNet
@@
0.D.4 Linear Probing on ImageNet
@@
0.D.5 SFT on ImageNet for SSL-pre-trained Models
@@
0.D.6 Segmentation for SSL-pre-trained Models
@@
0.D.7 Segmentation for Pyramid Transformers
We follow the setting of [12], which is almost the same as 0.D.6. We use a drop-path rate of 0.2 for the pyramid VisionLLaMA-B model.
0.D.8 Object Detection of Pyramid Transformers
We use the same setting as [12]. We use the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. All the models are trained for 36 epochs with a global batch size of 16. The initial learning rate is 1$\times$10-4 with 1000 iterations warm up and decayed by 10.0 at epoch 27 and 33. To avoid overfitting, we apply a $l_{2}$ weight decay for all models.
0.D.9 Object Detection of Plain Transformers
We use the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The training resolution is fixed as 1024 $\times$ 1024 as [38]. Our model is trained for 36 epochs with a global batch size of 64. The initial learning rate is 1$\times$10-4 with 1000 iterations warm up and decayed by 10.0 at epoch 27 and 33. We use a $L_{2}$ weight decay of 0.1. We also apply layer-wise learning rate decay with 0.7 as [38].
0.D.10 Image Generation of DiT-LLaMA
We use the same VAE as [52]. We make use of the AdamW optimizer with momentum $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We use a global batch size of 256 across all models. The learning rate is fixed as 1 $\times$ $10^{-4}$. The training resolution is 256 $\times$ 256. As for inference, we use 250 steps DDPM. We keep the default setting of ADM [21] without tuning. Specifically, we use $t_{\max}=1000$ linear schedule with $\beta$ from 0.0001 to 0.02 and learnable variance $\sigma_{\theta}$.
0.D.11 Image Generation of SiT-LLaMA
We use the same VAE as SD [52]. As for the ODE sampler, we utilize dopri5 and set ‘atol’ and ‘rtol’ to 1e66-6- 6 and 1e33-3- 3 respectively.
Appendix 0.E Architecture Setting
0.E.1 Pyramid VisionLLaMA
@@
0.E.2 Plain Transformer for Vision Understanding
@@
0.E.3 Plain Transformer for Generation
@@
Table 18: Architecture settings for VisionLLaMA on image generation.